What is the amount of energy required to increase the temperature of 150g of ice from -10°c to 0°c?
1
Expert's answer
2020-10-12T07:48:56-0400

We can find the required amount of energy as follows:


"Q = Q_1 + Q_2 = mc_{ice} \Delta T + mL_f,"

here, "Q_1" is the amount of energy required to change the temperature of the ice from "-10 \ ^{\circ}C" to "0 \ ^{\circ}C", "Q_2" is the amount of energy required to transform ice to water at "0 \ ^{\circ}C", "m = 0.15 \ kg" is the mass of ice, "c_{ice} = 2100 \ \dfrac{J}{kg \cdot \! ^{\circ}C}" is the specific heat capacity of ice, "L_f = 3.33 \cdot 10^5 \ \dfrac{J}{kg}" is the latent heat of fusion of ice and "\Delta T" is the change in temperature.

Then, we can calculate the amount of energy required to increase the temperature of 150g of ice from -10°C to 0°C:

"Q=0.15 \ kg \cdot 2100 \ \dfrac{J}{kg \cdot \! ^{\circ}C} \cdot 10 \ ^{\circ}C + 0.15 \ kg \cdot 3.33 \cdot 10^5 \ \dfrac{J}{kg} = 53100 \ J."

Answer:

"Q=53100 \ J."